Update 2013-10-17 00:30 UTC
Packet loss stopped, back to 100% connectivity. Awaiting RFO from network ops.
Update 2013-10-16 18:45 UTC
Service restored but 25-30% packet loss ongoing. Pending RFO (Reason for Outage) report from network ops.
Update 2013-10-16 16:45 UTC
Network ops are performing a full reload on the affected routers.
Update 2013-10-16 16:15 UTC
Network ops have identified the problem and are working to resolve it as soon as possible. Still no firm ETA available.
Update 2013-10-16 15:45 UTC
Our hardware is working and links are up, but no traffic is reaching our switches. Network ops are investigating.
2013-10-16 15:15 UTC
We are currently experiencing an outage at our Amsterdam 1 datacenter. Engineers are working on a solution but we do not yet have a firm ETA. Please check back shortly for an update.
Customers affected include shared Web hosting, reseller Web hosting, ns1.anu.net and ns2.anu.net DNS resolvers (ns3 is hosted in Chicago and is still up), and customers with virtual servers hosted on ams1-cloudmin.anu.net.